# Logarithmic Smoothing for Adaptive PAC-Bayesian Off-Policy Learning

### v1

@@ D.2 Proof of Proposition 1
Appendix E Proofs of Section 4
E.1 Proof of Theorem 4.1
@@ Appendix E Proofs of Section 4
Now for a given posterior $Q$ and a prior $P$, we apply the change of measure inequality [12, 13] for any $k$:
@@
Now; given a martingale is also a supermartingale, we can apply Ville’s inequality [14], which implies that with probability at least $1-\delta$, for any $k\geq 0$:
@@
Plugging this into Eq. 8 yields, with probability at least $1-\delta$, for any $k\geq 0$, any $Q$:
@@
E.2 Proof of Proposition 4.2
We need a preliminary proposition to prove Proposition 4.2.
@@
Proposition E.2.
For any adaptive algorithm Alg, data independent prior $P$, any $\boldsymbol{\lambda\in(0,1)},\delta\in(0,1]$, with probability at least $1-\delta$ over $S$, we have, for all $k\geq 0$, denoting by $\pi_{k+1}$ the output of Alg at time $k$, for all $Q\in\mathcal{P}(\Theta)$:
@@
with $M_{\lambda}(d_{\theta},\pi_{j})=\mathbb{E}_{(x,a,c)\sim P(\pi_{j})}\left[\frac{\left(\frac{d_{\theta}(a|x)}{\pi_{j}(a|x)}-1\right)^{2}c^{2}}{1-\lambda\left(\frac{d_{\theta}(a|x)}{\pi_{j}(a|x)}-1\right)c}\right].$ Now for a given posterior $Q$ and a prior $P$, we apply the change of measure inequality [12, 13] for any $k$:
@@
We can apply Ville’s inequality [14] which implies that with probability at least $1-\delta$, for any $k\geq 0$:
@@
Plugging this into Eq. 9 yields, with probability at least $1-\delta$, for any $k\geq 0$, any $Q$:
@@
Theorem E.4.
For any data independent prior $P$, any $\boldsymbol{\lambda\in(0,1)},\delta\in(0,1]$, with probability at least $1-\delta$, we have, denoting by $\pi_{j}$ the output of $\mathcal{A}$ at time $j$, for all $k\geq 0$:
@@
with $\pi_{k+1}$ defined as the output of the adaptive algorithm Algorithm 2 at time $k$.
@@
We know that by Thm. 4.1, for any data independent prior $P$, any $\boldsymbol{\lambda\in(0,1)},\delta\in(0,1]$, with probability at least $21-\delta/$ 2 over $S$, we have for all $k\geq 0$:
@@
Now, we exploit Proposition E.2, with probability at least $21-\delta/$ 2 over $S$, we have $\forall k\geq 0$:
@@
E.4 Proof of Theorem 4.4
We first start with a key intermediary result, derived from Lemma 4.3 coupled with Proposition 4.2.
Corollary E.5.
Let $\pi_{k+1}$ be the output of Algorithm 2 at time $k$ with parameter $\lambda\in(0,1)$. For any data independent prior $P$, any $\boldsymbol{\lambda\in(0,1)},\delta\in(0,1]$, with probability at least $1-\delta$, we have for all $k\geq 0$:
@@
Assume the result true for all $1\leq j\leq k$. By Cor. E.5, we have for any $\lambda\leq 1/2$:
@@
E.5 Proof of Corollary 4.5
Proof of Cor. 4.5.
@@
For $k\geq 0$, Assume the result true for all $j\leq k$. By Cor. E.5, we have:
@@
Appendix F Experimental Design and Additional Results
F.1 Detailed Setup

### v2

@@ B.2 Proof of Proposition 1
Appendix C SEQUENTIAL LEARNING ALGORITHMS WITH NON-IID CONTEXTS.
In the main text, we always considered an i.i.d. assumption on the contexts, which may be restrictive when considering real-life situations such as the behavior of an individual, susceptible to evolve through time. In what follows, we make no assumption on the context distribution and only assume that at time $k\in\{1\cdots,K\}$, conditionally to $\mathcal{F}_{k-1}$, $S_{k}=\left((x_{i},a_{i},c_{i})\right)_{i\in I_{k}}$ is constituted of mutually independent random variable. We show that we are still able to obtain a generalization bound, thanks to the supermartingale toolbox of Haddouche and Guedj [2022, 2023], Chugg et al. [2023]. We first start with some additional framework.
Framework
Assume that batches are of size $m$. We define, for all $k$, the risk at time $k$ of $\pi_{k}$ as
$R_{k}(\pi_{k})=\mathbb{E}_{x^{\prime}_{k}\sim\nu_{k},a^{\prime}_{k}\sim\pi_{k}(\cdot\mid x_{k})}[c(x^{\prime}_{k},a^{\prime}_{k})],$
where $\nu_{k}$ denotes, at time $k$, the conditional distribution of contexts with respect to the past (i.e. $\mathcal{F}_{k-1}$) and $x^{\prime}_{k},a^{\prime}_{k}$ are independent of $x_{k},a_{k}$. We also define, for any $\pi_{k},\pi_{k-1}$, the empirical risk at time $k$ as
$\hat{R}_{k}^{\lambda-LS}(\pi_{k},\pi_{k-1}):=\frac{1}{m}\sum_{i\in I_{k}}h_{LS}^{\lambda-Lin}(\pi_{k}(a_{i}\mid x_{i}),\pi_{k-1}(a_{i}\mid x_{i}),c_{i}),$
with $h_{LS}^{\lambda-Lin}$ defined in Sec.˜3. Given context distributions are evolving, we aim to learn a sequence of posteriors $Q_{k}$ which is $\mathcal{F}_{k}$- measurable from an initial prior $P=Q_{0}$ to make the cumulative risk $\frac{1}{K}\sum_{k=0}^{K}R_{k}(\pi_{Q_{k}})$ small.
Adapting our proof technique yields the following result.
Theorem C.1.
For any subbatch size $m>0$, any $\lambda>0,\delta\in(0,1]$, any learning algorithm $\mathcal{A}$ such that for all $k$, $Q_{k}=\mathcal{A}(S_{1},\cdots,S_{k})$ with probability at least $1-\delta$ over $S$, for any $K\geq 1$:
$\frac{1}{K+1}\sum_{k=0}^{K}R_{k}(\pi_{Q_{k}})\leq\frac{1}{K+1}\sum_{k=0}^{K}\hat{R}_{k}^{\lambda-LS}(\pi_{Q_{k}},\pi_{Q_{k-1}})+\frac{\sum_{k=0}^{K}\operatorname{KL}(Q_{k},Q_{k-1})+\log\left(\frac{1}{\delta}\right)}{\lambda(K+1)m}.$
This is dynamic counterpart of Sec.˜3.2, allowing shifting context distribution. Note that, similarly to Sec.˜3.2, this bound suggests a new learning algorithm detailed in Algorithm˜3. This algorithm only exploits $S_{k}$ and $\pi_{Q_{k-1}}$ to create $\pi_{Q_{k}}$, contrary to Algorithm˜1 which exploits all previous data $S_{0}\cup\cdots,\cup S_{k}$. The reason behind this discrepancy is that past data gathers outdated information as the context distribution may arbitrarily evolve. Then our theoretical result suggests to only consider current data. Note that the toolbox we use also allows an alternative result to Thm.˜4.1 for non-iid contexts, exploiting our novel regularized IPS.
1Input: Behavior policy $\pi_{Q_{-1}}$, a policy class $\mathcal{C}(\Theta)$, $\lambda>0$.
2 Initialize $P$ such that $\pi_{P}=\pi_{Q_{-1}}$ and $\mathcal{S}_{-1}=\{\}$.
3 for $k\geq 0$ do
4   Collect data $S_{k}$ of size $|I_{k}|=m$ with $\pi_{k}$.
5   Solve the optimization problem:
$\hat{Q}_{k+1}=\operatorname{argmin}_{Q\in\mathcal{C}(\Theta)}\left\{\hat{R}_{k}^{\lambda-LS}(\pi_{Q_{k}},\pi_{Q_{k-1}})+\frac{\operatorname{KL}(Q_{k},Q_{k-1})}{\lambda m}\right\}$
Set $\pi_{k+1}=\pi_{\hat{Q}_{k+1}}$.
Algorithm 3 Sequential Policy Learning via Logarithmic Smoothing for shifting context distributions
We then detail the proof of Thm.˜C.1.
Proof.
First, we follow the route described in Haddouche and Guedj [2023], Section 3.
For any $K>0$, we look for a $K$-tuple of probabilities. Thus, our predictor set of interest is $\mathcal{H}_{K}:=\mathcal{H}^{\otimes K}$ and then, our predictor $\theta^{K}$ is a tuple $(\theta_{1},..,\theta_{K})\in\mathcal{H}_{K}$.
Our goal is to apply the change of measure inequality on $\mathcal{H}_{K}$ to a specific function $f_{K}$. We define this function below, for any sample $S$ and any predictor $\theta^{K}=(\theta_{1},...,\theta_{K})$
$f_{K}(S,\theta^{K}):=\sum_{k=0}^{K}\underbrace{\sum_{i\in I_{k}}-\log\left(1-\lambda R_{k}(d_{\theta_{k}})\right)+\log\left(1-\lambda\frac{d_{\theta_{k}}(a_{i}\mid x_{i})c_{i}}{\pi_{Q_{k-1}}(a_{i}\mid x_{i})}\right)}_{:=Y_{k}(S,\theta_{k})}.$
Now for a given posterior tuple $Q_{1},...Q_{K}$ we define $Q^{K}=Q_{1}\otimes...\otimes Q_{K}$ and also $P^{K}=Q_{0}\otimes...\otimes Q_{K-1}$. We can now properly apply the change of measure inequality for any $m$:
$\sum_{k=0}^{K}\mathbb{E}_{\theta_{k}\sim Q_{k}}\left[Y_{k}(S,\theta_{k})\right]$ $=\mathbb{E}_{\theta^{k}\sim Q}\left[f_{K}(S,\theta^{K})\right]\leq\operatorname{KL}(Q^{K},P^{K})+\log\left(\mathbb{E}_{\theta^{K}\sim P^{K}}\exp(f_{K}(S,\theta^{K}))\right)$
$=\sum_{k=0}^{K}\operatorname{KL}(Q_{k},Q_{k-1})+\log\left(\mathbb{E}_{\theta^{K}\sim P^{K}}\exp(f_{K}(S,\theta^{K}))\right),$ (5)
where we define $f_{K}(S,\theta^{K}):=\sum_{k=0}^{K}Y_{k}(S,\theta_{k})$.
Lemma C.2.
The sequence $(M_{k}:=\mathbb{E}_{\theta^{K}\sim P^{K}}\exp(f_{K}(S,\theta^{K}))_{K\geq 0}$ is a non-negative martingale.
Proof.
We fix $K\geq 1$ and we recall that for any $k$, $Q_{k-1}$ is $\mathcal{F}_{k-1}$-measurable. We show that $\mathbb{E}_{K-1}[M_{K}]=M_{K-1}$.
$\mathbb{E}_{K-1}[M_{K}]$ $=\mathbb{E}_{K-1}\left[\mathbb{E}_{\theta^{K}\sim P^{K}}\exp(f_{K}(S,\theta^{K})\right]$
$=\mathbb{E}_{K-1}\left[\mathbb{E}_{\theta_{1},..,\theta_{K}\sim Q_{0}\otimes...\otimes Q_{K-1}}\exp(f_{K}(S,\theta^{K})\right]$
$=\mathbb{E}_{K-1}\left[\mathbb{E}_{\theta_{1},..,\theta_{K}\sim Q_{0}\otimes...\otimes Q_{K-1}}\left[\exp\left(f_{K-1}(S,\theta^{K-1})\right)\exp\left(\sum_{i\in I_{K}}\log\left(\frac{1-\lambda\frac{d_{\theta_{K}}(a_{i}\mid x_{i})c_{i}}{\pi_{Q_{K-1}}(a_{i}\mid x_{i})}}{1-\lambda R_{K}(d_{\theta_{K}})}\right)\right)\right]\right]$
$=M_{K-1}\mathbb{E}_{K-1}\left[\mathbb{E}_{\theta_{K}\sim Q_{K-1}}\left[\Pi_{i\in I_{K}}\frac{1-\lambda\frac{d_{\theta_{K}}(a_{i}\mid x_{i})c_{i}}{\pi_{Q_{K-1}}(a_{i}\mid x_{i})}}{1-\lambda R_{K}(d_{\theta_{K}})}\right]\right].$
The last line holding because $P^{K-1}=Q_{0}\otimes...\otimes Q_{K-2}$ is $\mathcal{F}_{K-1}$ measurable.
Now we exploit the fact that $Q_{K-1}$ is $\mathcal{F}_{K-1}$ measurable to apply a conditional Fubini lemma stated in Haddouche and Guedj [2022, Lemma D.3]. We have:
$\mathbb{E}_{K-1}\left[\mathbb{E}_{\theta_{K}\sim Q_{K-1}}\left[\Pi_{i\in I_{K}}\frac{1-\lambda\frac{d_{\theta_{K}}(a_{i}\mid x_{i})c_{i}}{\pi_{Q_{K-1}}(a_{i}\mid x_{i})}}{1-\lambda R_{K}(d_{\theta_{K}})}\right]\right]$ $=\mathbb{E}_{\theta_{K}\sim Q_{K-1}}\left[\mathbb{E}_{K-1}\left[\Pi_{i\in I_{K}}\frac{1-\lambda\frac{d_{\theta_{K}}(a_{i}\mid x_{i})c_{i}}{\pi_{Q_{K-1}}(a_{i}\mid x_{i})}}{1-\lambda R_{K}(d_{\theta_{K}})}\right]\right]$
$=\mathbb{E}_{\theta_{K}\sim Q_{K-1}}\left[\Pi_{i\in I_{K}}\frac{1-\lambda\mathbb{E}_{K-1}\left[\frac{d_{\theta_{K}}(a_{i}\mid x_{i})c_{i}}{\pi_{Q_{K-1}}(a_{i}\mid x_{i})}\right]}{1-\lambda R_{K}(d_{\theta_{K}})}\right].$
The last line holding as $R_{K}(d_{\theta_{K}})$ is $\mathcal{F}_{K-1}$-measurable for all $K,\theta_{K}$. This also holds as we exploited the assumption that, conditionally to $\mathcal{F}_{K-1}$, $S_{K}=\left((x_{i},a_{i},c_{i})\right)_{i\in I_{K}}$ is constituted of mutually independent random variable (thus an expectation of product equals a product of expectations). Now we remark that, for any $\theta_{K}$, $i\in I_{K}$:
$\mathbb{E}_{K-1}\left[\frac{d_{\theta_{K}}(a_{i}\mid x_{i})c_{i}}{\pi_{Q_{K-1}}(a_{i}\mid x_{i})}\right]$ $=\mathbb{E}_{x^{\prime}_{i}\sim\nu_{K}}\mathbb{E}_{a^{\prime}_{i}\sim\pi_{Q_{K-1}}(\cdot\mid x^{\prime}_{i})}\mathbb{E}_{c^{\prime}_{i}\sim p(\cdot\mid x^{\prime}_{i},a^{\prime}_{i})}\left[\frac{d_{\theta_{K}}(a^{\prime}_{i}\mid x^{\prime}_{i})c^{\prime}_{i}}{\pi_{Q_{K-1}}(a^{\prime}_{i}\mid x^{\prime}_{i})}\right]$
$=\mathbb{E}_{x^{\prime}_{i}\sim\nu_{K}}\mathbb{E}_{a^{\prime}_{i}\sim\pi_{Q_{K-1}}(\cdot\mid x^{\prime}_{i})}\left[\frac{d_{\theta_{K}}(a^{\prime}_{i}\mid x^{\prime}_{i})}{\pi_{Q_{K-1}}(a^{\prime}_{i}\mid x^{\prime}_{i})}\mathbb{E}_{c^{\prime}_{i}\sim p(\cdot\mid x^{\prime}_{i},a^{\prime}_{i})}\left(c^{\prime}_{i}\right)\right]$
$=\mathbb{E}_{x^{\prime}_{i}\sim\nu_{K}}\mathbb{E}_{a^{\prime}_{i}\sim\pi_{Q_{K-1}}(\cdot\mid x^{\prime}_{i})}\left[c\left(x^{\prime}_{i},d_{\theta_{K}}(x^{\prime}_{i})\right)\right]$
$=R_{K}(d_{\theta_{K}})$
Thus $\mathbb{E}_{K-1}[M_{K}]=M_{K-1}$, this concludes the lemma’s proof. ∎
$\mathbb{E}_{\theta^{k}\sim P^{k}}\exp(f_{k}(S,\theta))\leq\frac{1}{\delta}.$
Plugging this back into Appendix˜C yields, with probability at least $1-\delta$ over $S$, for any $K\geq 0$,
$\sum_{k=0}^{K}\mathbb{E}_{\theta_{k}\sim Q_{k}}\left[\sum_{i\in I_{k}}-\log\left(1-\lambda R_{k}(d_{\theta_{k}})\right)+\log\left(1-\lambda\frac{d_{\theta_{k}}(a_{i}\mid x_{i})c_{i}}{\pi_{Q_{k-1}}(a_{i}\mid x_{i})}\right)\right]\leq\sum_{k=0}^{K}\operatorname{KL}(Q_{k},Q_{k-1})+\log\left(\frac{1}{\delta}\right).$
Now remark that for any $k\geq 1,i\in I_{k},\theta_{k}$, because $d_{\theta_{k}}(a,x)=\mathds{1}\{f_{\theta_{k}}(x)=a\}\in\{0,1\}$, we have for any $(a,x)$:
$\log\left(1-\lambda\frac{d_{\theta_{k}}(a\mid x)c_{i}}{\pi_{Q_{k-1}}(a\mid x)}\right)=d_{\theta_{k}}(a,x)\log\left(1-\lambda\frac{c_{i}}{\pi_{Q_{k-1}}(a\mid x)}\right).$
We then have:
$\sum_{k=0}^{K}\mathbb{E}_{\theta_{k}\sim Q_{k}}\left[\sum_{i\in I_{k}}-\log\left(1-\lambda\frac{d_{\theta_{i}}(a_{i}\mid x_{i})c_{i}}{\pi_{i}^{0}(a_{i}\mid x_{i})}\right)\right]$ $=-\sum_{k=0}^{K}\sum_{i\in I_{k}}\pi_{Q_{k}}(a_{i}\mid x_{i})\log\left(1-\lambda\frac{c_{i}}{\pi_{Q_{k-1}}(a_{i}\mid x_{i})}\right)$
$=m\lambda\sum_{k=0}^{K}\hat{R}_{k}^{\lambda-LS}(\pi_{Q_{k}}).$
We also use twice the convexity of $x\rightarrow-\log(1-\lambda x)$ to obtain:
$\sum_{k=0}^{K}\mathbb{E}_{\theta_{k}\sim Q_{k}}\left[\sum_{i\in I_{k}}-\log\left(1-\lambda R_{k}(d_{\theta_{k}})\right)\right]$ $=m\sum_{k=0}^{K}\mathbb{E}_{\theta_{k}\sim Q_{k}}\left[-\log\left(1-\lambda R_{k}(d_{\theta_{k}})\right)\right]$
$\geq m\sum_{k=0}^{K}-\log\left(1-\lambda R_{k}(\pi_{Q_{k}})\right),$
$\geq-(K+1)m\log\left(1-\frac{\lambda}{K+1}\sum_{k=0}^{K}R_{k}(\pi_{Q_{k}})\right).$
Finally, mixing all those calculations together and dividing by $\lambda Km$ gives: with probability at least $1-\delta$, for any $K$:
$-\frac{1}{\lambda}\log\left(1-\frac{\lambda}{K+1}\sum_{k=0}^{K}R_{k}(\pi_{Q_{k}})\right)\leq\frac{1}{K+1}\sum_{k=0}^{K}\hat{R}_{k}^{\lambda-LS}(\pi_{Q_{k}},\pi_{Q_{k-1}})+\frac{\sum_{k=0}^{K}\operatorname{KL}(Q_{k},Q_{k-1})+\log\left(\frac{1}{\delta}\right)}{\lambda(K+1)m}$
Applying the function $\psi_{\lambda}$ Finally, mixing all those calculations together, dividing by $\lambda(k+1)m$ and applying the increasing function $\psi_{\lambda}(x):=\frac{1}{\lambda}(1-\exp(-\lambda x)$ gives:
$\frac{1}{K+1}\sum_{k=0}^{K}R_{k}(\pi_{Q_{k}})\leq\psi_{\lambda}\left(\frac{1}{K+1}\sum_{k=0}^{K}\hat{R}_{k}^{\lambda-LS}(\pi_{Q_{k}},\pi_{Q_{k-1}})+\frac{\sum_{k=0}^{K}\operatorname{KL}(Q_{k},Q_{k-1})+\log\left(\frac{1}{\delta}\right)}{\lambda(K+1)m}\right).$
Finally, using that $\psi_{\lambda}(x)\leq x$ concludes the proof. ∎
Appendix D PROOFS OF SECTION 4
D.1 Proof of Theorem 4.1
@@ Appendix D PROOFS OF SECTION 4
Now for a given posterior $Q$ and a prior $P$, we apply the change of measure inequality [Csiszár, 1975, Donsker and Varadhan, 1976] for any $k$:
@@
Now; given a martingale is also a supermartingale, we can apply Ville’s inequality [Doob, 1939], which implies that with probability at least $1-\delta$, for any $k\geq 0$:
@@
Plugging this into Eq.˜6 yields, with probability at least $1-\delta$, for any $k\geq 0$, any $Q$:
@@
D.2 Proof of Proposition 4.2
We need a preliminary proposition to prove Proposition˜4.2.
@@
Proposition D.2.
For any sequential algorithm Alg, data independent prior $P$, any $\boldsymbol{\lambda\in(0,1)},\delta\in(0,1]$, with probability at least $1-\delta$ over $S$, we have, for all $k\geq 0$, denoting by $\pi_{k+1}$ the output of Alg at time $k$, for all $Q\in\mathcal{P}(\Theta)$:
@@
with $M_{\lambda}(d_{\theta},\pi_{j})=\mathbb{E}_{(x,a,c)\sim P(\pi_{j})}\left[\frac{\left(\frac{d_{\theta}(a|x)}{\pi_{j}(a|x)}-1\right)^{2}c^{2}}{1-\lambda\left(\frac{d_{\theta}(a|x)}{\pi_{j}(a|x)}-1\right)c}\right].$ Now for a given posterior $Q$ and a prior $P$, we apply the change of measure inequality Csiszár [1975], Donsker and Varadhan [1976] for any $k$:
@@
We can apply Ville’s inequality Doob [1939] which implies that with probability at least $1-\delta$, for any $k\geq 0$:
@@
Plugging this into Eq.˜7 yields, with probability at least $1-\delta$, for any $k\geq 0$, any $Q$:
@@
Theorem D.4.
For any data independent prior $P$, any $\boldsymbol{\lambda\in(0,1)},\delta\in(0,1]$, with probability at least $1-\delta$, we have, denoting by $\pi_{j}$ the output of Algorithm˜2 at time $j$, for all $k\geq 0$:
@@
with $\pi_{k+1}$ defined as the output of the sequential algorithm Algorithm˜2 at time $k$.
@@
We know that by Thm.˜4.1, for any data independent prior $P$, any $\boldsymbol{\lambda\in(0,1)},\delta\in(0,1]$, with probability at least $1-\delta/2$ over $S$, we have for all $k\geq 0$:
@@
Now, we exploit Proposition˜D.2, with probability at least $1-\delta/2$ over $S$, we have $\forall k\geq 0$:
@@
D.4 Proof of Theorem 4.4
We first start with a key intermediary result, derived from Lemma 4.3 coupled with Proposition˜4.2.
Corollary D.5.
Let $\pi_{k+1}$ be the output of Algorithm˜2 at time $k$ with parameter $\lambda\in(0,1)$. For any data independent prior $P$, any $\boldsymbol{\lambda\in(0,1)},\delta\in(0,1]$, with probability at least $1-\delta$, we have for all $k\geq 0$:
@@
Assume the result true for all $1\leq j\leq k$. By Cor.˜D.5, we have for any $\lambda\leq 1/2$:
@@
D.5 Proof of Corollary 4.5
Proof of Cor.˜4.5.
@@
For $k\geq 0$, Assume the result true for all $j\leq k$. By Cor.˜D.5, we have:
@@
Appendix E EXPERIMENTAL DESIGN AND ADDITIONAL RESULTS
E.1 Detailed Setup